Wakefield Industries purchased a delivery vehicle on January 1, 2019, for $30,000. They paid sales taxes of $2,000 and one year’s insurance premium of $2,000 on the same day. The company estimated a residual value of $3,000 at the end of the 5 year expected service life.

Expected usage (in miles) of the new delivery vehicle
2019 16,000
2020 22,000
2021 25,000
2022 21,000
2023 16,000

Mileage estimates were based on expected increases in customers and the addition of a second delivery vehicle in 2022.

What was depreciation expense of the vehicle in 2021 under the sum-of-the-years’-digits and double-declining balance depreciation methods, respectively?

Select one:

a. $5,800: $4,608

b. $6,200; $4,176

c. $6,400; $4,176

d. $6,200; $4,896

e. $6,800; $4,896
Answer #1

Cost of Van = 30000+2000+2000 = 34000

Expected miles = 16000+22000+25000+21000+16000 = 100000

Double decline rate = 100/5*2 = 40%

2021 Dep under sum of year digit method = (34000-3000)*3/15 = 6200

2021 Dep under Double decline rate = 34000*60%*60%*40% = 4896

So answer is d) $6200; $4896